40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 214 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  6204   Tue Jan 17 02:44:59 2012 kiwamuUpdateCDSawg not working

AWG is not working. This needs to be fixed.

I could set the channel and the parameters in the AWGGUI screen, but it never inject signals to the realtime system.

  6203   Tue Jan 17 02:27:49 2012 kiwamuUpdateLSCfringe tests : all the suspensions are innocent

I did a quick test to check a hypothesis that PRM is interfering with some other optics in the single bounce configuration.

I shook all the suspensions (except the MC mirrors) at 3 Hz in POS, PIT and YAW with an amplitude of 1000 counts.

No effects were found in the REFL demod signals.

So it is NOT a fringe effect caused by the other suspended mirrors.

Quote from #6202

The REFL11 and REFL55 demod signals show high frequency noise depending on how big signals go to the POS actuator of PRM.

Is PRM making some fringes with some other optics ??   

 

  6202   Tue Jan 17 01:02:07 2012 kiwamuUpdateLSCglitch hunting in REFL RFPDs : strange

A very strange thing is going on.

The REFL11 and REFL55 demod signals show high frequency noise depending on how big signals go to the POS actuator of PRM.

This noise shows up even when the beam is single-bounced back from PRM ( the rest of the suspensions are misaligned) and it's very repeatable.

Any idea ?? Am I crazy ?? Is PRM making some fringes with some other optics ??

 

 

(background)

 The most annoying thing in the central part locking is glitches showing up in the LSC error signals (#6183).
The symptom is that when the motion in PRCL at 3 Hz becomes louder, somehow we get glitches in both the MICH and PRCL error signals.
In the frequency domain, those glitches are mostly contribute to a frequency band of about 30 - 100 Hz.
Last Thursday Koji and I locked the half PRM (PRMI with either ITMX or ITMY misaligned) to see if we still have the glitches in this simpler configuration.
Indeed there were the same kind of glitches --a loud 3 Hz motion triggers the glitches.
It was shown particularly in the REFL11 signal but not so much in the REFL33 while AS55 didn't show any glitches.
 
 
(Still glitches even in the single bounce beam)
   We were suspecting some kind of coupling from a beam jitter, so that the 3 Hz motion somehow brings the beam spot to a bad place somewhere in the REFL paths.
I misaligned all the suspensions except for PRM such that the beam directly bounces back from PRM and go to the REFL port.
Indeed there still were glitches in the REFL11 and REFL55 demod signals. It showed up once per 30 sec or so and pushes up the noise floor around 30 - 100 Hz.
There might be a little bit of glitches also in the REFL33, but the ADC noise floor and the expected glitch noise level were comparable and hence it was difficult to see the glitches in REFL33.
 

(Glitch is related to the PRM POS actuation)

 In the single-bounce configuration I started shaking the PIT and YAW motions of PRM at 3 Hz using the realtime LOCKIN oscillator to see if I can reproduce the glitches.
However no significant glitches were found in this test.
Then I started shaking the POS instead of the angular DOFs, and found that it causes the glitches.
At this point it didn't look like a glitch any more, it became more like a stationary noise.
 
 The attached screen shot is the noise spectrum of the REFL11_I.
The red curve is the one taken when I injected the 3 Hz excitation in POS by the LOCKIN oscillator.
The excitation is at 3 Hz with an amplitude of 1000 counts.
As a comparison I plotted the same spectrum when no excitation was injected and it is plotted in pink.
 

 PRMsingle_bounce.png

It seems there is a cut off frequency at 100 Hz.
This frequency depends on the amplitude of the excitation -- increasing the amplitude brings the cut off frequency higher.
This noise spectrum didn't change with and without the oplevs and local damping.
 

(Possible scenario)

A possible reason that I can think of right now is : PRM is interfering with some other optics for some reason.
But if it's true, why I didn't see any fringes in the AS demod signals in the half PRM configuration ?
 

Quote from #6183

 We tried to figure out what is causing spikes in the REFL33 signal, which is used to lock PRCL.

No useful information was obtained tonight and it is still under investigation.

  6201   Sun Jan 15 12:18:00 2012 DenUpdateAdaptive Filteringrunning time

 In order to figure out what downsampling ratio we can take, we need to determine the running time of the fxlms_filter() function. If the filter length is equal to 5000, downsampling ratio is equal to 1, number of witness channels is 1 then with ordinary compilation without speed optimization one call runs for 0.054 ms (milli seconds). The test was done on the 3 GHz Intel processor. With speed optimization flags the situation is better

-O1 0.014 ms, -O3 0.012 ms

However, Alex said that speed optimization is not supported at RCG because it produce unstable execution time for some reason. However, by default the kernel should optimize for size -Os. With this flag the running time is also 0.012 ms. We should check if the front-end machine compilers indeed use -Os flag during the compilation and also play with speed optimization flags. Flags -O3 and -Os together might also give some speed improvement.

But for now we have time value = 0.012 ms as running time for 5000 coefficient filter, 1 witness channel and downsample ratio = 1. Now, we need to check how this time is scaled if we change the parameters.

5000 cofficients - 0.012 ms

10000 coefficients - 0.024 ms

15000 coefficients - 0.036 ms

20000 coefficients - 0.048 ms

We can see that filter length scaling is linear. Now we check downsampling ratio

ratio=1 - 0.048 ms

ratio=2 - 0.024 ms

ratio=4 - 0.012 ms

Running time on the dependance of downsample ratio is also linear as well as on the dependence of the number of witness channels and degrees of freedom. 

If we want to filter 8 DOF with approximately 10 witness channels for each DOF, then 5000 length filter will make 1 cycle for ~1 ms, that is good enough to make the downsample ratio equal to 4.

Things get a little bit complicated when the code is called for the first time. Some time is taken to initialize variables and check input data. As a result the running time of the first cycle is ~0.1 ms for 1 DOF that is ~10 times more then running time of an ordinary cycle. This effect takes place at this moment when one presses reset button in the c1oaf model - the filter becomes suspended for a while. To solve this problem the initialization should be divided by several (~10) parts.

  6200   Sun Jan 15 11:40:30 2012 DenUpdateAdaptive Filteringdownsampling

Here for the downsampling process we use a low-pass Bessel digital filter of order 6, normalized cut-off frequency = 0.1. In the plot presented below we compare the results with downsampling ratio = 1, 2, 4.

2048_downsampling.png

We can see that increasing the downsampling ratio, we increase the error of the filter. Moreover, the error at some particular frequency f seems to depend on the ratio f/Fs, where Fs - sampling frequency (2048 Hz) devided by downsampling ratio. Error is the same for all curves below 1 Hz but then begins to increase as we increase the sownsampling ratio. In order to figure out what the problem is - mistake in the filter code, inaccurate upsample algorithm or this is NLMS particularity, I've changed sampling frequency in the chans/daq/C1PEM.ini and C1IOO.ini files from 2048 Hz to 512 Hz for corresponding channels. Now, we compare the error from the filter working with 2048 Hz frequency, downsampling ratio = 4, low-pass filter = Bessel of order 6, normalized cut-off = 0.1 and filter working with 512 Hz sampling frequency, without downsampling and with corresponding Bessel low-pass filter with normalized frequency 0.4.

2048_512.png

MC_F measurement at 2048 Hz was done during the day, for that reason red curved is slightly higher then green in the resonance frequencies. But still we can see that these two cases are very much alike. For this reason, it seems that NLMS filter works better with higher sampling frequencies.

  6199   Sun Jan 15 10:28:02 2012 DenUpdateAdaptive Filteringdelays

We can account for delays in the oaf system by compensating it in the adaptive path of the filter. But using only this procedure is not enough. Parameters mu and tau should be chosen accurately:

w = (1 - tau) * w;

w += mu * dw / norm;

NLMS algorithm without considering delays works well for mode cleaner length and gur1 seismometer signals, significantly reducing MC_F  with parameters mu=1, tau=0. These parameters are considered because nlms algorithm should converge with the highest speed when mu=1. However, if the system has a delay so at time moment n:

error_signal [n] = desired_signal [n] - filter_output [n-delay];

then the adaptive filter diverges for the same parameters mu=1 and tau=0 even for delay=1. For that reason we make the same calculations with tau = 1e-4 and tau = 1e-2 without reducing mu conserving the adaptation rate and get the same result as nlms algorithm without delays. Next figure shows MC_F signal, error after applying e-nlms filter with tau=1e-4 and tau=1e-2. "e-" is added to show that  a small number (epsilon) is added to the norm of the signal in order to prevent the filter from diverging in the beginning of the process when the norm is not well-determined yet.

2048_tau.png

The test was done offline with the sampling frequency 2048 Hz, without downsampling and any filters. We can see that tau=1e-4 is still not enough, tau=1e-3 or tau=1e-2 is as good as nlms without delays, tau=1e-1 and high are also bad.

Correctly choosing tau we have some freedom for delay compensation in the adaptation path. This is important as we do not know exactly what is the delay in the real system. We can measure it approximately. In order to figure out the range of reasonable delay errors we make a test with delay = 1, but to the adaptation path we give delays from 0 to 10. It turns out that adaptation path delays greater then 5 make the filter diverge, delays in the range 0-3 produce a reasonable error. In the figure below errors with adaptation path delays = 1 (correct) and 3 are presented.

2048_delays.png

  6198   Sat Jan 14 00:50:08 2012 rana, kojiConfigurationIOOTowards coating thermal noise measurement with RefCav / MC beat

Koji asked aloud tonight if we could measure the coating thermal noise of the refcav optics by beating the refcav light with the MC_TRANS light. Then we looked at our calculations for the noises:

Displacement noise of T=200ppm silica/tantala coating on a 1" silica substrate with a 300 micron beam spot = 1e-18 * sqrt(100 Hz / f) m/rHz.

Displacement noise from coating thermal in the MC is roughly smaller by the beam size ratio (1.8 mm / 0.3 mm). Some differences due to 3 mirrors and more layers on MC2 than the others, but those are small factors.

So, the frequency noise from the refcav should by larger than the MC thermal noise by a total factor of (1.8 / 0.3) * (13 m / 8 inches) ~ 400.

Another way to say it is that the effective strain noise in the RC is (1e-18 / 0.200) = 5e-18 /rHz. This translates into (5e-18 * 13) = 6.5e-17 m/rHz in the MC. (in frequency noise its 1.5 mHz/rHz).

I have measured the frequency noise in the LLO MC to be at this level back in 2009, so it seems possible to use our RC + MC to measure coating thermal noise by the length amplification factor and compete with Frank+Tara.

 

So today we set up the Jenny RC temperature setup to lock the LWE NPRO to the RC and then set up the beat note with the IFO REFL beam on the AS table. By using the 2 laser beat, we are avoiding the VCO phase noise issue which used to limit the PSL frequency noise at ~0.01 Hz/rHz. To do this we have reworked some of the optics on the PSL and AS tables, but I think its been done without disturbing the beams for the regular locking. Beat note has been found, but the NPRO has still not been locked to the RC - next we setup the lockin amp, dither the PZT, and then use the New Focus lock box to lock it to the RC.

You might think that its hard to measure this since the MC has ~1 MHz frequency fluctuations and we want to measure down to 1e-4 Hz. But, in fact, we can just use a 200 m MFD with a LT1128 preamp. Then we use the MFD to stabilize the MC length to the refcav and just use the control + error signal of the MFD setup as the coating thermal noise measurement.

 

Note: Beat found at ~40deg for the aux laser. The aux laser is on but the shutter is closed.
The AS camera seems to be hosed. Need a bit of alignment. (KA) ==> Fixed. (Jan 15)

  6197   Fri Jan 13 17:40:38 2012 ZachUpdateRF Systemfoam box and temp controller taken off of PSL table

I stole the foam EOM box and the temperature controller circuit from the PSL table so that I could continue with the RAM measurements here at the ATF.

That is all.

  6196   Fri Jan 13 16:16:05 2012 Leo SingerUpdateStewart platformFlexure type for leg joints

I had been thinking of using this flexure for the bearings for the leg joints, but I finally realized that it was not the right type of bearing.  The joints for the Stewart platform need to be free to both yaw and pitch, but this bearing actually constrains yaw (while leaving out-of-plane translation free).

  6195   Fri Jan 13 00:51:40 2012 Leo SingerUpdateStewart platformFrequency-dependent requirements for Stewart platform

Below are revised design parameters for the Stewart platform based on ground motion measurements.

The goal is that the actuators should be able to exceed ground motion by a healthy factor (say, two decades in amplitude) across a range from about .1 Hz to 500 Hz.  I would like to stitch together data from at least two seismometers, an accelerometer, and (if one is available) a microphone, but since today this week I was only able to retrieve data from one of the Guralps, I will use just that for now.

The spectra below, spanning GPS times 1010311450--1010321450, show the x, y, and z axes of one of the Guralps.  Since the Guralp's sensitivity cuts off at 50 Hz or so, I assumed that the ground velocity continues to fall as f-1, but eventually flattens at acoustic frequencies.  The black line shows a very coarse, visual, piecewise linear fit to these spectra.  The corner frequencies are at 0.1, 0.4, 10, 100, and 500 Hz.  From 0.1 to 0.4 Hz, the dependence is f-2, covering the upper edge of what I presume is the microseismic peak.  From 0.4 to 10 Hz, the fit is flat at 2x10-7 m/s/sqrt(Hz).  Then, the fit is f-1 up to 100 Hz.  Finally, the fit remains flat out to 500 Hz.

ground_velocity_spectrum.png

Outside this band of interest, I chose the velocity requirement based on practical considerations.  At high frequencies, the force requirement should go to zero, so the velocity requirement should go as f--2 or faster at high frequencies.  At low frequencies, the displacement requirement should be finite, so the velocity requirement should go as f or faster.

The figure below shows the velocity spectrum extended to DC and infinite frequency using these considerations, and the derived acceleration and displacement requirements.

requirements.png

As a starting point for the design of the platform and the selection of the actuators, let's assume a payload of ~12 kg.  Let's multiply this by 1.5 as a guess for the additional mass of the top platform itself, to make 18 kg.  For the acceleration, let's take the maximum value at any frequency for the acceleration requirement, ~6x10-5 m/s2, which occurs at 500 Hz.  From my previous investigations, I know that for the optimal Stewart platform geometry the actuator force requirement is (2+sqrt(3))/(3 sqrt(2))~0.88 of the net force requirement.  Finally, let's throw in as factor of 100 so that the platform beats ground motion by a factor of 100.  Altogether, the actuator force requirement, which is always of the same order of magnitude as the force requirement, is

(12)(1.5)(6x10-5)(0.88)(100) ~ 10 mN.

Next, the torque requirement.  According to <http://www.iris.edu/hq/instrumentation_meeting/files/pdfs/rotation_iris_igel.pdf>, for a plane shear wave traveling in a medium with phase velocity c, the acceleration a(x, t) is related to the angular rate W'(x, t) through

a(x, t) / W'(x, t) = -2 c.

This implies that |W''(f)| = |a(f)| pi f / c,

where W''(f) is the amplitude spectral density of the angular acceleration and a(f) of the transverse linear acceleration.  I assume that the medium is cement, which according to Wolfram Alpha has a shear modulus of mu = 2.2 GPa and about the density of water: rho ~ 1000 kg/m3.  The shear wave speed in concrete is c = sqrt(mu / rho) ~ 1500 m/s.

The maximum of the acceleration requirement graph is, again, 6x10-5 m/s2 at 500 Hz..  According to Janeen's SolidWorks drawing, the largest principal moment of inertia of the SOS is about 0.26 kg m2.  Including the same fudge factor of (1.5)(100), the net torque requirement is

(0.26) (1.5) (6x10-5) (100) pi (500) / (1500) N m ~ 2.5x10-3 N m.

The quotient of the torque and force requirements is about 0.25 m, so, using some of my previous results, the dimensions of the platform should be as follows:

radius of top plate = 0.25 m,

radius of bottom plate = 2 * 0.25 m = 0.5 m, and

plate separation in home position = sqrt(3) * 0.25 m = 0.43 m.

 

One last thing: perhaps the static load should be taken up directly by the piezos.  If this is the case, then we might rather take the force requirement as being

(10 m/s2)(1.5)(12 kg) = 180 N.

An actuator that can exert a dynamic force of 180 N would easily meet the ground motion requirements by a huge margin.  The dimensions of the platform could also be reduced.  The alternative, I suppose, would be for each piezo to be mechanically in parallel with some sort of passive component to take up some of the static load.

  6194   Thu Jan 12 23:19:56 2012 KojiUpdateSUSc1iscex is fine now

c1iscex is working as before and the optic is damped.


What I checked

1. I went to the X-end rack. I found the io-chassis was turned off.

2. I shutdown c1iscex, turned off, and turned on everything. Again, we did not have any signal from the ADC into c1scx model.

However, I found that c1x01 indicates healthy ADC signals.
This means that the connection between the IOP and the c1scx model was wrong ==> Simulated Plant

3. Burtrestored X'mas eve snapshot. This restored the gains and matrices as well as C1:SCX-SIM_SWITCH
which switches the input between the real ADCs and simulated plant.

4. The signals came back to c1scx.

  6193   Thu Jan 12 23:13:42 2012 KojiConfigurationWIKI-40M UpdateUnable to create Wiki page

Quote:

 I can't create a new page on the 40m wiki.  The page that I was trying to create is

http://blue.ligo-wa.caltech.edu:8000/40m/Stewart

I get this message when I try to save the new page:

Page could not get locked. Unexpected error (errno=13).

This address for wiki is obsolete. Recently it was switched to https://wiki-40m.ligo.caltech.edu/
Jamie is working on automatic redirection from the old wiki to the new place.

The new one uses albert.einstein authentication.

 

  6192   Thu Jan 12 21:22:16 2012 Leo SingerConfigurationWIKI-40M UpdateUnable to create Wiki page

 I can't create a new page on the 40m wiki.  The page that I was trying to create is

http://blue.ligo-wa.caltech.edu:8000/40m/Stewart

I get this message when I try to save the new page:

Page could not get locked. Unexpected error (errno=13).

  6191   Thu Jan 12 11:08:23 2012 Leo SingerUpdatePEMFunky spectrum from STS-2

I am trying to stitch together spectra from seismometers and accelerometers to produce a ground motion spectrum from Hz to 100's of Hz.  I was able to retrieve data from two seismometers, GUR1 and STS_1, but not from any of the accelerometers.  The GUR1 spectrum is qualitatively similar to other plots that I have seen, but the STS_1 spectrum looks strange: the X axis spectrum is falling off as ~1/f, but the Y and Z spectra are pretty flat.  All three axes have a few lines that they may share in common and that they may share with GUR1.

See attached plot.

Attachment 1: spectrum.jpg
spectrum.jpg
  6190   Thu Jan 12 10:11:59 2012 steveUpdatePEMoptical table top with 1" wall

Quote:

3/4 " thick colored acrylic material will be used in this air tight design. Surgical tubing for o-ring. We may have to put an o-ring into the bottom to have it really air tight.

Feedtrouhs: www.roxtec.com

The top drawing is not ready. It will have handle and industrial grade L-handle lock pin to hold cover down. There will be 2  one inch od post in the midle of the table to hold the cover and lock  the ball pin.

I'm waiting for your inputs, so I can send this preliminary design out for quote.

 

 The present plan to go with clear cast acrylic plexiglass 1" thick side wall  and  two  clear  1/2 thick top cover.

The inside would be lined with light braun YAG safety window sheets 0.14"  VLT ~25%  OD 4 @ 532nm & OD 5 @ 1064nm

 

Attachment 1: 101122012atc.PDF
101122012atc.PDF
  6188   Thu Jan 12 07:44:21 2012 steveUpdateSUSsapphire wire standoff quote


On Tue, Dec 20, 2011 at 11:37 PM, Rana Adhikari wrote:
Steve,

We are thinking of replacing a few metal parts in the suspension with sapphire/ruby. We want to replace the standoff (which is Al) with sapphire and also the top plate where the wire is clamped.

For the top, we should make a cutout for a little plate in the existing crossbar. In the cutout would go a little mounting plate. Then the clamping plate (which is now steel) we would also replace with a Sapphire one.

I don't think it matters whether its sapphire or ruby, just has to be very hard.

Can you please get some quotes on the parts?

rana

Attachment 1: standoff.PDF
standoff.PDF
  6187   Thu Jan 12 03:05:02 2012 kiwamuUpdateLSCOSA installed in AS

[John / Valera / Kiwamu]

 We installed a new weapon, an optical spectrum analyzer in the AS port.

Like we used to do in the old days, two BNC cables were newly laid down and they bring the output of the OSA to the control room to monitor the spectrum with an oscilloscope.

 

(Some notes)

The photo diode of the OSA was replaced by a Thorlab PDA100A to amplify the signals.

The carrier peak is at about 6.9 V and the f1 and f2 sidebands peaks are at about 40 mV when the beam is in straight shot (everything is misaligned except ITMY and BS).

According to a rough calculation, those numbers correspond to a modulation depth of about 0.16 or so.

The depth agree with what Mirko measured before (#5519)

 

  6186   Wed Jan 11 21:35:33 2012 Max De JongUpdateGeneralProjector

I finished mounting the new projector. The projector and computer monitor now display information.

  6185   Wed Jan 11 14:06:28 2012 Leo SingerHowToComputer Scripts / ProgramsHowTo for getting data from NDS off site

 This may or may not be general knowledge already, but Jamie and I added a HowTo explaining how to retrieve channel data from the frame builder via NDS, but off site on one's own computer.  See the Wiki page:

https://wiki-40m.ligo.caltech.edu/How_To/NDS

  6184   Tue Jan 10 09:17:23 2012 steveUpdatePhotosstrawman's visiters
Attachment 1: P1080491.JPG
P1080491.JPG
Attachment 2: P1080492.JPG
P1080492.JPG
  6183   Tue Jan 10 00:09:33 2012 kiwamuUpdateLSCspike hunting in REFL33

[John / Kiwamu]

 We tried to figure out what is causing spikes in the REFL33 signal, which is used to lock PRCL.

No useful information was obtained tonight and it is still under investigation.

 


(Background)

 One thing preventing us from doing smooth measurements of the noise budget and the sensing matrix is some sharp spikes in the LSC error signals.

For example when we lock PRMI with REFL33 and AS55 fedback to PRCL and MICH respectively, both the REFL33 and AS55 signals show some spikes in time series.

Those spikes then bring the noise spectra higher than how they should be.

So for the reason, taking the noise budget doesn't give us much information about the interferometer rather than there are spikes.

Also the sensing matrix measurement has been suffered from those spikes, which excite the impulse responses of the low pass filters in the LOCKIN detection systems a lot.

 

(What we did)

 We looked into the actual analog signals to see if there are indeed spikes or not before they are acquired to the ADCs.

But we didn't find any corresponding spikes in the signals that are after the mixers.

It maybe because the signals we looked into didn't have high enough SNR because they were coming out from the monitor lemo outputs on the demod boards.

 Then we thought the spikes are from the whitening circuits, due to some kind of saturation.

We decreased the gain of the whitening filters by a factor of 10, but it didn't help and the spikes were still there.

  6182   Mon Jan 9 23:52:15 2012 kiwamuUpdateCDSSUS channels not accessible from dataviewer

[John / Kiwamu]

 We found that some of the suspensions channels (for example C1:SUS-BS_POS_IN1 and etc) were not accessible from dataviewer for some reasons.

So far it seems none of the channels associated with c1sus are accessible from dataviewer.

  6181   Mon Jan 9 13:19:09 2012 kiwamuUpdateSUSETMX damping restored

No we can't do that because the c1scx model is not working properly.

If you look into the real time controller screen you will find what I mean.

Quote from #6180

ETMX sus damping restored

  6180   Mon Jan 9 08:47:35 2012 steveUpdateSUSETMX damping restored

ETMX sus damping restored

  6179   Fri Jan 6 20:10:48 2012 ranaUpdateCDSframebuilder back online, using NTP time syncronization

 

 You (Jamie) should talk with Rolf and Alex to find out what framebuilder they will support for > 3 years. Then we should buy that along with the adapter card which allows us to use the same RAID we now have for the frames.

  6178   Fri Jan 6 19:17:07 2012 Max De JongUpdateGeneralMounted projector

I mounted the new projector to the pipe where the old projector was attached. The mounting hardware wasn't designed for attaching to a pipe, but with Steve's help I mounted the projector. The projector is angled away from the area of the wall designated as the screen, and I am going to meet with Steve Monday to fix this.

  6177   Fri Jan 6 14:31:54 2012 JamieUpdateCDSframebuilder back online, using NTP time syncronization

The framebuilder is back online now, minus it's symmetricom GPS card.  The card seems to have failed entirely, and was not able to be made to work at downs either.  It has been entirely removed from fb.

As a fall back, the system has been made to work off of the system NTP-based time synchronization.  The latest symmetricom driver, which is part of the RCG 2.4 branch, will fall back to using local time if the GPS synchronization fails.  The new driver was compiled from our local checkout of the 2.4 source in the new to-be-used-in-the-future rtscore directory:

controls@fb ~ 0$ diff {/opt/rtcds/rtscore/branches/branch-2.4/src/drv/symmetricom,/lib/modules/2.6.34.1/kernel/drivers/symmetricom}/symmetricom.ko
controls@fb ~ 0$ 

The driver was reloaded.   daqd was also linked against the last running stable version and restarted:

controls@fb ~ 0$ ls -al $(readlink -f /opt/rtcds/caltech/c1/target/fb/daqd)
-rwxr-xr-x 1 controls controls 6592694 Dec 15 21:09 /opt/rtcds/caltech/c1/target/fb/daqd.20120104
controls@fb ~ 0$ 

We'll have to keep an eye on the system, to see that it continues to record data properly, and that the fb and the front-ends remain in sync.

The question now is what do we do moving forward.  CDS is not supporting the symmetricom cards anymore, and have moved to using Spectracom GPS/IRIG-B cards.  However, Downs has neither at the moment.  Even if we get a new Spectracom card, it might not work in this older Sun hardware, in which case we might need to consider upgrading the framebuilder to a new machine (one supported by CDS).

  6176   Fri Jan 6 11:49:13 2012 JamieUpdateCDSframebuilder taken offline to diagnose problem with symmetricom timing card

Alex and I have taken the framebuilder offline to try to see what's wrong with the symmetricom card.  We have removed the card from the chassis and Alex has taken it back to downs to do some more debugging.

We have been formulating some alternate methods to get timing to the fb in case we can't end up getting the card working.

  6175   Fri Jan 6 01:00:56 2012 kiwamuUpdateCDSc1scx out of sync

Both the c1scx and its IOP realtime processes became out of sync.

Initially I found that the c1scx didn't show any ADC signals, though the sync sign was green.

Then I software-rebooted the c1iscex machine and then it became out of sync.

For tonight this is fine because I am concentrating on the central part anyway.

  6174   Thu Jan 5 20:40:21 2012 JamieUpdateCDSRTS upgrade aborted; restored to previous settings; fb symmetricom card failing?

After running into more problems with the upgrade, I eventually decided to abort todays upgrade attempt, and revert back to where we were this morning (RTS 2.1).  I'll try to follow this with a fuller report explaining what problems I encountered when attempting the upgrade.

However, when Alex and I were trying to figure out what was going wrong in the upgrade, it appears that the fb symmetricom card lost the ability to sync with the GPS receiver.  When the symmeticom module is loaded, dmesg shows the following:

[  285.591880] Symmetricom GPS card on bus 6; device 0
[  285.591887] PIC BASE 2 address = fc1ff800
[  285.591924] Remapped 0x17e2800
[  285.591932] Current time 947125171s 94264us 800ns 
[  285.591940] Current time 947125171s 94272us 600ns 
[  285.591947] Current time 947125171s 94280us 200ns 
[  285.591955] Current time 947125171s 94287us 700ns 
[  285.591963] Current time 947125171s 94295us 800ns 
[  285.591970] Current time 947125171s 94303us 300ns 
[  285.591978] Current time 947125171s 94310us 800ns 
[  285.591985] Current time 947125171s 94318us 300ns 
[  285.591993] Current time 947125171s 94325us 800ns 
[  285.592001] Current time 947125171s 94333us 900ns 
[  285.592005] Flywheeling, unlocked...

Because of this, the daqd doesn't get the proper timing signal, and consequently is out of sync with the timing from the models.

It's completely unclear what caused this to happen.  The card seemed to be working all day today, then Alex and I were trying to debug some other(maybe?) timing issues and the symmetricom card all of a sudden stopped syncing to the GPS.  We tried rebooting the frame builder and even tried pulling all the power to the machine, but it never came back up.  We checked the GPS signal itself and to the extend that we know what that signal is supposed to look like it looked ok.

I speculate that this is also the cause of the problems were were seeing earlier in the week.  Maybe the symmetricom card has just been acting flaky, and something we did pushed it over the edge.

Anyway, we will try to replace it tomorrow, but Alex is skeptical that we have a replacement of this same card.  There may be a newer Spectracom card we can use, but there may be problems using it on the old sun hardware that the fb is currently running on.  We'll see.

In the mean time, the daqd is running rogue, off of it's own timing.  Surprisingly all of the models are currently showing 0x0 status, which means no problems.  It doesn't seem to be recording any data, though.  Hopefully we'll get it all sorted out tomorrow.

  6173   Thu Jan 5 09:59:27 2012 JamieUpdateCDSRTS/RCG/DAQ UPGRADE TO COMMENCE

RTS/RCG/DAQ UPGRADE TO COMMENCE

I will be attempting (again) to upgrade the RTS, including the RCG and the daqd, to version 2.4 today.  The RTS will be offline until further notice.

  6172   Thu Jan 5 09:08:48 2012 steveUpdateGeneralframebuilder is back

Dataviewer is recovered after heavy new year party.

Attachment 1: dataviewerisback.png
dataviewerisback.png
  6171   Wed Jan 4 16:40:52 2012 JamieUpdateComputersfront-end fb communication restored

Communication between the front end models and the framebuilder has been restored.  I'm not sure exactly what the issue was, but rebuilding the framebuilder daqd executable and restarting seems to have fixed the issue.

I suspect that the problem might have had to do with how I left things after the last attempt to upgrade to RCG 2.4.  Maybe the daqd that was running was linked against some library that I accidentally moved after starting the daqd process.  It would have kept running fine as was, but if the process died and was attempted to be started again, it's broken linking might have kept it from running correctly.  I don't have any other explanation.

It turns out this was not (best I can tell) related to the new year time sync issues that wer seen at the sites.

  6170   Wed Jan 4 16:22:30 2012 kiwamuUpdateLSCpower normalization in LSC : modification done

The dynamic power normalization system has been modified such that the normalization happen after the LSC input matrix.

The attached screen shot below tells you how the signals flow.
The red circled region in the picture is the place where the power normalization are performed.
pow_norm.png
 
The dynamic normalization will be activated once you put some numbers into the elements in the matrix.
Otherwise the error signals are always normalized by 1.

Quote from #6158

It turned out that the power normalization need a modification.

I will work on it tomorrow and it will take approximately 2 hours to finish the modification.

 

  6169   Wed Jan 4 11:47:08 2012 JamieSummaryComputer Scripts / Programsmedm directory clean-up

Quote:

 I moved the following old and obsolete MEDM directories to an archive /cvs/cds/rtcds/caltech/c1/medm

  • c1vga
  • c1vgl
  • c1gpt
  • c1gpv
  • c1nio
  • c1spy
  • c1spx

None of these models were present in /cvs/cds/rtcds/caltech/c1/target

Remember that we don't use /cvs/cds/rtcds anymore.  Everything should be in /opt/rtcds now.

  6168   Wed Jan 4 09:06:50 2012 steveUpdateComputerspossible front-end timing issue

Quote:

Quote:

Is there a reason the framebuilder status light is red for all the front ends?

Also, I reenabled PRM watchdog.

Apparently there is a bug in the timing cards having to do with the new year roll-over that is causing front-end problems.  From Rolf:

For systems using the Spectracom IRIG-B cards for timing information, the code did not properly roll over the time for
2012 (still thinks it is 2011 and get reports from DAQ of timing errors (0x4000)). I have made a temporary fix for this
in the controller.c code in branch-2.3, branch-2.4 and release 2.3.1. 

I was going to check to see if the 40m is suffering from this. I'll be over to see if that's the problem.

 The problem is the same as yesterday.

Attachment 1: rtntstat.png
rtntstat.png
  6167   Wed Jan 4 05:02:58 2012 kiwamuUpdateLSCSidebands measurement at POP
Just a quick report:
I did the first attempt to measure the recycling gains of the sidebands in the DRMI configuration (sidebands resonant condition)
by looking at the output of the POP22/110 RFPD.
Because this time what I measured is some absolute values of the sidebands power,
it doesn't tell us anything quantitatively until we calibrate it or compare it with similar data.
So I need to measure the same things in some different configurations (e.g. PRMI, SRMI, etc.)
in order to extract some useful information from the measurement.
 
The attached picture is the display of a power spectrum analyzer looking at the output of the POP22/110 broadband RFPD
while the DRMI (in the sideband resonant condition) was kept locked.
You can see that 111 MHz (twice of 55 MHz) is prominent. Also there are several peaks at 11, 22, 44 and 66 MHz.
SB_DRMI.png
  6166   Wed Jan 4 03:03:24 2012 kiwamuUpdateLSClocking activity tonight and beyond
Last night and tonight, I was doing a kind of rehabilitation -- locking PRMI and DRMI with the new trigger system.
Although MC wasn't so awesome (#6164), I confirmed that the DRMI can stay locked with the conventional RFPD combination (#4760).
Additionally I have modified the IFO configure scripts, such that they also automatically restore the thresholds values for triggering.
The scripts are available in the C1IFO_CONFIGURE screen as usual.

 

       Locking plan            

Here is a plan in my mind and these are basically the details of the gantt chart (#6143):

  • (1 day task) Measurement of the recycling gains of the RF sidebands with the PRMI and DRMI configuration, using POP22/110 RFPD.
    • I need to have confidence that I am really locking the DRMI with SRC resonating to 55 MHz.
    • Also those values will enable us to estimate losses and mode matching again (maybe ?).
  • (3-4 days task) Measurement of the sensing matrix using the multiple-LOCKIN system.
    • Write a script to automatically measure the sensing matrix. This must be easy.
      • The results will enable us to diagonalize the input matrix and therefore it eventually gives more solid lock of the DRMI
      • Also it will give us the optical gains of 3f signals. So this is actually a step toward the 3f signal check.
  • (3-4 days task) Noise budgeting on the 3f signals
    • This is a very important part of the DRMI characterization because the results will tell us whether we can hold the DRMI lock with a sufficient SNR or not.
    • If it turns out that they don't have good SNRs, we then have to come up with some ideas to improve the SNRs.
  • (Extra fun task depending on schedule) 3f DRMI lock + Y arm ALS
    • If the beat-box electronics are not available by the time when the work above are completed, I will do this fun task.
    • Probably it is better to start preparing the common mode servo electronics because it will be needed anyway.

 

  6165   Wed Jan 4 02:43:40 2012 ranaUpdateGeneralActuators for Stewart platform

Quote:

 I checked on the two single-axis shakers that are present at the 40m that Steve pointed out:

  • Brüel & Kjær type 4809, rated for 45 N peak, and
  • Brüel & Kjær type 4810, rated for 10 N peak.

Neither of these meet the force requirement of 2.04 kN peak.

 Time to lower your expectations!

Do we really need 40 microns at 500 Hz? Or perhaps should there be a frequency dependent displacement requirement?

  6164   Wed Jan 4 00:43:06 2012 kiwamuUpdateIOOMC became flaky

I don't know what exactly is going on, but MC became flaky and it's been frequently unlocked.

I have turned off the MC WFS servo to check if the WFSs are doing something bad. But it still tends to be unlocked without the WFS servo.

Right now it doesn't stay locked for more than 10 min.

  6163   Tue Jan 3 20:42:05 2012 Leo SingerUpdateGeneralActuators for Stewart platform

 I checked on the two single-axis shakers that are present at the 40m that Steve pointed out:

  • Brüel & Kjær type 4809, rated for 45 N peak, and
  • Brüel & Kjær type 4810, rated for 10 N peak.

Neither of these meet the force requirement of 2.04 kN peak.

  6162   Tue Jan 3 18:40:08 2012 AidanSummaryComputer Scripts / Programsmedm directory clean-up

 I moved the following old and obsolete MEDM directories to an archive /cvs/cds/rtcds/caltech/c1/medm

  • c1vga
  • c1vgl
  • c1gpt
  • c1gpv
  • c1nio
  • c1spy
  • c1spx

None of these models were present in /cvs/cds/rtcds/caltech/c1/target

  6161   Tue Jan 3 17:46:38 2012 Updated Stewart platform design requirementsUpdateGeneralUpdated design requirements for a Stewart platform

I updated the design requirements for the Stewart platform.  I weighed the unloaded "dirty" SOS that was sitting on the workbench in the control room; its mass is 11 kg.  Steve suggested that the OSEMs (not installed on this model) would add another 0.5 kg.  From the specs in the final SOS design document, LIGO-T970135, I added 0.25 kg for the optic itself; I am therefore taking the total payload mass to be 11.75 kg.  (Now, the upper stage of the Stewart platform itself will likely add a nontrivial amount, but I am not worrying about this yet.)

I have e-mailed Janeen Romie to obtain the actual center of mass and principal moments of inertia of the platform.  I also cooked up a simple scheme to measure both quantities, should this information not be available.  It would involve rigidly mounting the dirty SOS to a rigid bar hung from a pivot.  By translating the mount point in two dimensions and measuring the period of the pendulum, I ought to be able to find the center of mass and moments of inertia by multilinear regression.  However, this elaborate scheme is not necessary to just compute some ballpark figures; it could wait until a later stage in the design.  For the time being, I just rescaled the moment of inertia proportional to the increase in mass, such that the torque-to-force ratio is unchanged.

As such, the design requirements are now 

  • linear travel: 40 microns peak to peak (based on SOS design requirements in LIGO-T950011)
  • angular travel: 3 mrad peak to peak (based on SOS design requirements in LIGO-T950011)
  • payload mass: 11.75 kg (measured unloaded mass, plus educated guesses about combined mass of OSEMs and optic)
  • payload moment of inertia: 0.0232 kg m^2 (wild guess)
  • bandwidth: 500 Hz (suggestion of Rana and Koji: ~kHz)

From these assumptions, the revised actuator requirements and dimensions are:

  • peak actuator force: 2.04 kN
  • minimum radius of top platform: 15 cm
  • minimum radius of bottom platform: 30 cm
  • minimum height: 26 cm

See the attached PDF document.

It appears that the actuator that I had originally nominated, PI's model P-225.80, would very nearly meet the actuator force requirement.  Steve also pointed out the following single-axis shakers that are already in use in the 40m:

  • Brüel & Kjær type 4809
  • Brüel & Kjær type 4810

I want to find out if either of these would meet the present need, but I'm waiting on a response from the manufacturer to get access to the data sheets.

 

Attachment 1: stewart.pdf
stewart.pdf stewart.pdf stewart.pdf stewart.pdf stewart.pdf stewart.pdf stewart.pdf stewart.pdf
  6160   Tue Jan 3 17:25:27 2012 KojiUpdatePSLFound the laser was off

I found the PSL laser has been off for four hours. Nobody seemed to know why.

I just turned it on and it is now providing about 10% lower power compared with one before the shutdown.
Let's keep the eyes on the power if it can recover as the housing gets warm.

  6159   Tue Jan 3 15:49:27 2012 JamieUpdateComputerspossible front-end timing issue

Quote:

Is there a reason the framebuilder status light is red for all the front ends?

Also, I reenabled PRM watchdog.

Apparently there is a bug in the timing cards having to do with the new year roll-over that is causing front-end problems.  From Rolf:

For systems using the Spectracom IRIG-B cards for timing information, the code did not properly roll over the time for
2012 (still thinks it is 2011 and get reports from DAQ of timing errors (0x4000)). I have made a temporary fix for this
in the controller.c code in branch-2.3, branch-2.4 and release 2.3.1. 

I was going to check to see if the 40m is suffering from this. I'll be over to see if that's the problem.

  6158   Tue Jan 3 15:48:39 2012 kiwamuUpdateLSCpower normalization in LSC

It turned out that the power normalization need a modification.

I will work on it tomorrow and it will take approximately 2 hours to finish the modification.

 

     Concept of Power Normalization         

Koji pointed out that the dynamic power normalization, which I have installed(#6156),  should be placed after the LSC input matrix rather than before the matrix.
Now let us review the concept of the power normalization to avoid some confusions.
We will need two kinds of power normalizations as follows:
  1.  Static power normalization, which should be placed before the input matrix.
  2.  Dynamic power normalization, which should be placed after the input matrix.
 The static power normalization will be applied to each I and Q signals in all the LSC signals and also DCPD signals.
This normalization is supposed to cancel the effects from the incident laser power and depths of the phase modulations.
Because the variations in the laser power and modulation depth are expected to be relatively slow, we will apply static normalizations.
 
 The dynamic power normalization will be applied to the DOFs error signals, for example C1:LSC-DARM_IN and so on.
This normalization is supposed to cancel the effect of the internal states of the interferometer, for example alignments.
In addition to it, this dynamic normalization can expand the linear range of the error signals.

Quote from #6156

Now a power normalization is doable for the LSC error signals.

 

  6157   Tue Jan 3 15:45:04 2012 JenneUpdateComputersFB?

Is there a reason the framebuilder status light is red for all the front ends?

Also, I reenabled PRM watchdog.

  6156   Fri Dec 30 22:05:16 2011 kiwamuUpdateLSCpower normalization in LSC

Now a power normalization is doable for the LSC error signals.

It is working fine, but at some point we may want to have some kind of a saturation filter or limiter to avoid dividing a signal by a small number.

 

 (How to set the normalization)

  •   Click a small matrix panel on the LSC OVERVIEW window (shown in the attached screen shot below).
    •     This will give you a pop-up-window, which shows a matrix to route the normalization signals
POW_NORM_MTRX.png
  •   Choose a numerator channel, which you want to divide, and choose denominator channels, which you want to use as a power normalization factor.
  •   Put some number in the corresponding matrix elements.
  •   Once you put a non-zero element in the matrix, the corresponding numerator channel will be divided by the specified denominator channels.
    •     Otherwise the static normalization factors (e.g. C1:LSC-AS55_POW_NORM, etc.,) will be used for the denominator.
  6155   Fri Dec 30 02:16:48 2011 kiwamuUpdateGreen LockingYarm ALS : high frequency noise reduced

The high frequency noise, which has been a dominant noise above 30 Hz in the Y arm ALS (#6133), decreased by a factor of 5.

This reduction was done by increasing the modulation depth at the Y end PDH locking. Now the noise floor at 100 Hz went to 0.2 pm/sqrtHz.

However the noise source is not yet identified and hence it needs a further investigation.

 

 The attached figure is the sensor noises, which were taken from the beat-note signal while the arm was locked by the IR-PDH.
The orange curve is the one before I changed the modulation depth and the red curve is the one taken after I increased the modulation depth.
The high frequency noise went down from 1 pm/sqrt Hz to 0.2 pm/sqr tHz at 100 Hz.
 
Yarm_ALS_2011Dec29.png

 (Increasing the modulation depth)

  Actually I was going to check the RAM noise at the Y end PDH locking as I planed (#6143).
During some preparation for it, I found that there had been a 20 dB attenuator in the modulation LO path.
The reason we have kept it is that somehow a big modulation depth made the reflected DC light noisier.
For curiosity I removed it to see what will happen and took the noise spectra. Then the noise decreased as shown in the plot above.
It means the noise source was like a kind of sensor noise, whose level depends on the responsivity of the sensor.
As far as I can tell, it is not the dark noise or shot noise according to some quick measurements.
  6154   Wed Dec 28 14:13:16 2011 kiwamuUpdateGreen LockingALS feedback on MC2

I added an ALS feedback path on the MC2 suspension and this path will enable us to stabilise the MC length using the ALS scheme.

  The actual digital signal is transmitted from the c1gcv realtime controller to the c1mcs realtime controller through the c1rfm realtime process.
Or in terms of the machines, the signal is transmitted from C1IOO to C1SUS via the reflective memory network.
 
The attached figure is a screen shot of the MC2 position controller screen.  The new ALS path is emphasized by a purple circle in the figure.
MC2_ALS.png

Quote from #5888

Leaving a note on the ALS feedback before I forget:

The MC2 suspension needs to have an input for the ALS feedback in the realtime model like ETMs.

 

ELOG V3.1.3-